# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

### v1

@@ Appendix B Experiments
B.1. Baseline LLMs
To demonstrate the robustness of our proposed datasets, we evaluate eight fast-thinking LLMs and four slow-thinking reasoning LLMs on our benchmark tasks. We leverage vLLM (Kwon et al., 2023) to run all open-source LLMs locally, utilizing two Nvidia H100 GPUs. For the proprietary LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to reduce inference cost and perform inference in batches. In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1. Below is the list of LLMs used in our experiments:
Fast-thinking LLMs:
  • Mistral-7B-Instruct-v0.3 Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) is an instruct fine-tuned version of the Mistral-7B-v0.3 which exhibits extended vocabulary and supports the v3 tokenizer.
  • Qwen2.5-7B-Instruct Qwen2.5-7B-Instruct (Team, 2024a) is one of the advanced Qwen series LLMs. It exhibits better instruction following and long text generating capability.
  • LLaMa-3.1-8B-Instruct LLaMa-3.1-8B-Instruct (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
  • LLaMa-3.2-3B-Instruct LLaMa-3.2-3B-Instruct (Grattafiori et al., 2024) is a lightweight 3B version of LLaMA-3.2, prioritizing the rapid response and the deployment on low-resource devices.
  • Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct (Team, 2024a) is one of the latest LLMs from Qwen and serves as the biggest Qwen-2.5 series LLM with 72B parameters.
  • LLaMa-3.3-70B-Instruct LLaMa-3.3-70B-Instruct (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
  • GPT-4o GPT-4o (OpenAI, 2024b) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing  [^5] To assess its superiority, we leverage the most recent GPT-4o-2024-11-20 and GPT-4o-mini-2024-07-18 (at the time point when we did this work) as the benchmark evaluator, respectively..
  • GPT-4o-mini gpt-4o-mini (OpenAI, 2024a) is a compact variant of the GPT-4o (OpenAI, 2024b), which is optimized for faster inference and cost efficiency 5.
Slow-thinking reasoning LLMs:
  • DeepSeek-R1-Distill-Llama-8B DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025) is fine-tuned from the open-source Llama-3.1-8B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
  • DeepSeek-R1-Distill-Qwen-32B DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025) denotes the distilled version from the well-trained Qwen2.5-32B without any Reinforce Learning stage.
  • QWQ-32B-Preview QWQ-32B-Preview (Team, 2024b) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
  • GPT-o3-mini GPT-o3-mini (OpenAI, 2025) is designed with a focus on enhancing LLMs’ reasoning capabilities. It leverages the Chain of Thought (CoT) to break down complex problems into several simpler steps to achieve this objective.
The retrievers in our experiments are built using LlamaIndex (Liu, 2022). We set the number of retrieved chunks $u$ and its size $l$ to be $3$ and $512$ respectively.
B.2. TTP Temporal Position Analysis
We conduct a detailed breakdown of the performance in four indicative LLMs based on each ATT&CK tactic associated to the questions in AttackSeqBench. Based on Table 10, the mean and standard deviation (SD) for each tactic and between each benchmark task is shown. We observe that the overall highest standard deviation occurs within the AttackSeq-Tactic task for all models, indicating that it is challenging for LLMs to correctly answer questions in AttackSeq-Tactic.
Table 10. Results of performance of four representative models on all the 14 tactics, and the associative statistics of mean and standard deviation (SD).
@@ B.2. TTP Temporal Position Analysis
B.3. Error Analysis
To better understand the limitations of LLMs in analyzing attack sequences, we examine a specific cases where GPT-4o successfully and fails to answer the given question correctly (refer to Table 11). In the first example, the model incorrectly associates the technique in answer choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic. While this technique involves the execution of malicious payloads based on the ATT&CK KB, it is associated to only the following tactics, Persistence, Privilege Escalation and Defense Evasion. This incorrect interpretation suggests that the model struggles in distinguishing the inherent ambiguity found in TTP descriptions, which affects their ability to analyze attack sequences.
@@ B.3. Error Analysis
Table 11. Error analysis of GPT-4o in AttackSeq-Technique. The LLM incorrectly associates technique T1574 with Execution tactic
Question & Answer Choices LLM Response
@@
Question: Before establishing persistence through creating a Run key in the registry, which ATT&CK technique most likely facilitated the execution of the backdoor FinalBot.exe after the WMI discovery process identified the presence of antivirus software? Answer Choices: A) T1574-Hijack Execution Flow B) T1003.001-LSASS Memory C) T1059.008-Network Device CLI D) T1129-Shared Modules (Correct Answer) **A: T1574-Hijack Execution Flow** This technique involves hijacking the execution flow of a legitimate process to execute malicious code. DLL sideloading is a common method for this technique. The CTI report explicitly mentions that Sidewinder used DLL sideloading (‘cmpbk32.dll‘ to execute ‘Srvstr.exe‘), which aligns with this technique. This likely facilitated the execution of the backdoor ‘FinalBot.exe‘ after the WMI discovery process. **B: T1003.001-LSASS Memory**: This technique involves dumping credentials from the LSASS process memory. There is no mention in the CTI report of credential dumping or any activity related to LSASS. This choice is irrelevant. **C: T1059.008-Network Device CLI**: This technique involves using a command-line interface (CLI) to interact with network devices. The CTI report does not mention any interaction with network devices or the use of a CLI. This choice is irrelevant. **D: T1129-Shared Modules** This technique involves executing malicious code by loading shared modules. While DLL sideloading involves shared modules, the specific ATT&CK technique for DLL sideloading is T1574-Hijack Execution Flow. This choice is less precise than option A. The most plausible sequence of TTPs involves the use of **T1574-Hijack Execution Flow** (via DLL sideloading) to execute the backdoor ‘FinalBot.exe‘ after the WMI discovery process and before establishing persistence (Wrong Conclusion).
Appendix C Prompts Strategy
C.1. Question Generation Prompts
@@
We utilize few-shot prompting for each of our benchmark tasks, the prompts for AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No is shown in Tables 12, 13, 14, 15 respectively.
Table 12. Few-shot prompt for question generation for AttackSeq-Tactic. We insturct the LLM to generate a question based on the given CTI outline, where the answer to the question is the given ATT&CK TTP.
@@
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. [End of Inputs]
@@
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, where the answer to the question is the given MITRE ATT&CK tactic. The question should focus on inferring the given tactic by using the attack sequence based on the remaining tactics in the CTI report. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is the given MITRE ATT&CK tactic. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, which ATT&CK tactic most likely occurred before Ke3chang achieved persistence by adding a Run key? Answer: Execution - Question: Which ATT&CK tactic most likely occurred before Axiom gained initial access to the victim’s network using SQL injection? Answer: Resource Development - Question: Which ATT&CK tactic most likely occurred after Ke3chang establishes connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Exfiltration - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, which ATT&CK tactic most likely occurred before APT41 used Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Execution [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
@@
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
Table 13. Few-shot prompt for question generation for AttackSeq-Technique.
@@
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. 3. A **MITRE ATT&CK Technique** present in the CTI report. [End of Inputs]
@@
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic and technique given below.
Table 14. Few-shot prompt for question generation for AttackSeq-Procedure-Yes.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive four parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. 3. A **MITRE ATT&CK Technique** present in the CTI report. 4. 4. A list of **Procedures** present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet. [End of Inputs]
@@
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.

### v2

@@ Appendix B Experiments
B.1 TTP Temporal Position Analysis
To better understand the LLMs’ capabilities in attack sequence analysis, we conduct fine-grained analysis based on each stage within the attack sequences of MITRE ATT&CK®. We illustrate the performance of two LLMs and two LRMs across all benchmark tasks in the Regular setting in Figure 9 and show the corresponding values, the mean and standard deviation (SD) of these LLMs and LRMs on each tactic and benchmark task in Table 6. We identify four overachieving attack phases to categorize the ATT&CK tactics in attack sequences: (1) Initial Intrusion Phase; (2) Exploitation Phase; (3) Stealth Expansion Phase; (4) Objective Orchestration Phase. It is worth noting that our categorization follows Tactics, as each Technique and Procedure in MITRE ATT&CK® is uniquely mapped to a specific Tactic within a given attack sequence.
Figure 9: The performance comparison between two LLMs and two LRMs on each tactic in each benchmark task.
Table 6: Results of performance of two LLMs and two LRMs on all the 14 tactics, and the corresponding statistics of mean and standard deviation.
@@ B.1 TTP Temporal Position Analysis
We summarize our findings into three key points: (1) Comparing the performance of three benchmark tasks, the overall highest standard deviation occurs within the AttackSeq-Tactic task for all LLMs and LRMs, suggesting that the LLMs’ and LRMs’ performance has the largest variance in this taskand further highlighting the challenge of inferring tactics within attack sequences; (2) In contrast, the standard deviation for AttackSeq-Procedure is the lowest for all LLMs, suggesting more consistency in the LLMs’ and LRMs’ performance among the different attack sequence stages; (3) GPT-4o consistently outperforms other LLMs and LRMs across most benchmark tasks, particularly in the Exploitation Phase and the Stealth Expansion Phase. However, the remaining LLMs achieve comparable performance in the remaining two phases, especially in tactics at the start and end of the attack sequences (e.g., Reconnaissance and Impact). This suggests that TTPs within rarely observed tactics are more challenging for the LLM and LRMs to interpret.
B.2 Case Study
To better understand the superiority and limitation of the advanced LLMs in analyzing attack sequences, we run GPT-4o multiple times to collect both successful and failed responses to the same questions under the identical experimental setting, which are presented in Table 7. For the Wrong case, GPT-4o incorrectly associates the technique of choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic. Although this technique indeed involves the execution of malicious payloads based on the ATT&CK knowledge base, it is only associated to the following three tactics: Persistence, Privilege Escalation, and Defense Evasion. This incorrect interpretation suggests that GPT-4o struggles in distinguishing the inherent ambiguity found in TTP descriptions, thereby affecting their ability to analyze attack sequences. Regarding the Right one, GPT-4o correctly identifies T1129-Shared Modules as the most plausible technique, which belongs to the Execution tactic  [^8] https://attack.mitre.org/techniques/T1129/ and serves as the executable files that are loaded into processes to provide access to execute malicious payloads. By selecting this option, GPT-4o demonstrates its ability to reason over the attack sequence: after WMI discovery detects the presence of antivirus, shared modules would facilitate the execution of “FinalBot.exe” to bypass basic defenses. This correct interpretation is beneficial to effectively link the ambiguous textual cues with the appropriate tactic/technique entities, thereby improving its reliability in analyzing attack sequences.
@@ B.2 Case Study
Table 7: Error analysis of GPT-4o in AttackSeq-Technique, where GPT-4o incorrectly associates technique T1574-Hijack Execution Flow with the Execution tactic.
Q & A Cases LLM Response
@@
Question: Before establishing persistence through creating a Run key in the registry, which ATT&CK technique most likely facilitated the execution of the backdoor FinalBot.exe after the WMI discovery process identified the presence of antivirus software? Options: A) T1574-Hijack Execution Flow B) T1003.001-LSASS Memory C) T1059.008-Network Device CLI D) T1129-Shared Modules (Correct Answer) Wrong A: T1574-Hijack Execution Flow This technique involves hijacking the execution flow of a legitimate process to execute malicious code. DLL sideloading is a common method for this technique. The CTI report explicitly mentions that Sidewinder used DLL sideloading (‘cmpbk32.dll‘ to execute ‘Srvstr.exe‘), which aligns with this technique. This likely facilitated the execution of the backdoor ‘FinalBot.exe‘ after the WMI discovery process. B: T1003.001-LSASS Memory: This technique involves dumping credentials from the LSASS process memory. There is no mention in the CTI report of credential dumping or any activity related to LSASS. This choice is irrelevant. C: T1059.008-Network Device CLI: This technique involves using a command-line interface (CLI) to interact with network devices. The CTI report does not mention any interaction with network devices or the use of a CLI. This choice is irrelevant. D: T1129-Shared Modules: This technique involves executing malicious code by loading shared modules. While DLL sideloading involves shared modules, the specific ATT&CK technique for DLL sideloading is T1574-Hijack Execution Flow. This choice is less precise than option A. Thus, the most plausible sequence of TTPs involves the use of T1574-Hijack Execution Flow (via DLL sideloading) to execute the backdoor ‘FinalBot.exe‘ after the WMI discovery process and before establishing persistence. (Wrong Conclusion)
Right A: T1574-Hijack Execution Flow: This technique involves exploiting or redirecting the flow of execution to execute untrusted code. While this could potentially execute malware, it doesn’t directly relate to the sequence of WMI discovery leading to the execution of ‘FinalBot.exe‘. B: T1003.001-LSASS Memory: This technique involves accessing credentials in memory by dumping the LSASS process. This is unrelated to executing a backdoor like ‘FinalBot.exe‘ after WMI discovery. C: T1059.008-Network Device CLI: This technique involves command-line interfaces for network devices, which is unrelated to executing programs on a system post-discovery of antivirus software. D: T1129-Shared Modules: This technique involves executing malicious payloads using shared libraries or modules. If ‘FinalBot.exe‘ were executed via shared modules, this process could occur after determining system defenses, making it relevant to the scenario. Answer: D: T1129-Shared Modules. This technique likely facilitated the execution of ‘FinalBot.exe‘ after WMI discovery detected antivirus presence, allowing for a subtle execution method bypassing basic defenses. (Correct Conclusion)
Table 8: Performance comparison between three embedding models (abbreviate as Emb. M.).
#Emb. M. BGE OPENAI ATT&CK-BERT
Tasks Tactic Technique Procedure Tactic Technique Procedure Tactic Technique Procedure
Llama-3.1-8B 0.4751 0.5974 0.5243 0.4838 0.6103 0.5435 0.4820 0.5980 0.5334
GPT-4o 0.5522 0.6860 0.6319 0.5616 0.6578 0.6482 0.5687 0.7016 0.6216
R1 (Llama-8B) 0.4905 0.5740 0.5226 0.4932 0.5696 0.5150 0.4651 0.5804 0.5104
GPT o3-mini 0.5115 0.5853 0.6474 0.5192 0.5827 0.6474 0.5245 0.5874 0.6414
B.3 Impact of Embedding Models within RAG-empowered Setting
The semantics within CTI reports contain a large volume of domain-specific technical terminologies, where the accuracy of retrievers in identifying the most relevant tactics, techniques and procedures critically influences LLMs’ performance in RAG-empowered settings. To further examine this issue and mitigate the potential knowledge bias introduced by BGE-EN-ICL, we incorporate two additional embedding models into our RAG-empowered setting, namely OpenAI’s text-embedding-3-large (OpenAI, 2024b) and a domain-adapted ATT&CK BERT  (Abdeen et al., 2023b) fine-tuned on the specific cybersecurity data. We conduct performance comparisons between two representative LLMs and two LRMs based on these above embedding models in Table 8. We observe that the three embedding models exhibit comparable performance across different benchmark tasks and settings, with only marginal differences. Among them, BGE-EN-ICL proves to be the most cost-efficient, generalizable, and effective choice, and thus we primarily report model performance based on this embedding throughout the paper. Notably, although ATT&CK-BERT contains far fewer parameters than BGE-EN-ICL (110M vs. 7B), it achieves comparable performance, underscoring the importance of injecting domain-specific security knowledge into LLMs and pointing to a promising direction for future work.
Appendix C Prompt Templates
C.1 Question Generation Prompt Templates
@@
We utilize few-shot prompt templates for question generation of each of our benchmark tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No), the corresponding prompts are shown in Box C.1, Box C.1, Box C.1, Box C.1 respectively.
Box 1: Few-shot prompt for question generation for AttackSeq-Tactic. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]:
You will receive two parts:
@@
1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, where the answer to the question is the given MITRE ATT&CK tactic. The question should focus on inferring the given tactic by using the attack sequence based on the remaining tactics in the CTI report. [Steps]:
Please follow these steps:
@@
1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is the given MITRE ATT&CK tactic. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]:
Example Questions:
- Question: After gaining initial access through compromised VPN accounts, which ATT&CK tactic most likely occurred before Ke3chang achieved persistence by adding a Run key? Answer: Execution - Question: Which ATT&CK tactic most likely occurred before Axiom gained initial access to the victim’s network using SQL injection? Answer: Resource Development - Question: Which ATT&CK tactic most likely occurred after Ke3chang establishes connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Exfiltration - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, which ATT&CK tactic most likely occurred before APT41 used Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Execution [End of Examples]
3. Provide the Question-Answer Pair:
@@
- Please follow the output format: “Question: <insert question here> Answer: <insert answer here> [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
@@
Box 2: Few-shot prompt for question generation for AttackSeq-Technique. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. 3. A MITRE ATT&CK Technique present in the CTI report. [End of Inputs]
@@
3. Provide the Question-Answer Pair:
- Please follow the output format: “Question: <insert question here> Answer: <insert answer here>”. [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic and technique given below.
Box 3: Few-shot prompt for question generation for AttackSeq-Procedure-Yes. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive four parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. 3. A MITRE ATT&CK Technique present in the CTI report. 4. 4. A list of Procedures present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet. [End of Inputs]
@@
3. Provide the Question-Answer Pair: